# Learning the Basis: A Kolmogorov-Arnold Network Approach Embedding Green’s Function Priors

### v2

@@ Abstract
The Method of Moments (MoM) is constrained by the usage of static, geometry-defined basis functions, such as the Rao-Wilton-Glisson (RWG) basis. This letter reframes electromagnetic modeling around a learnable basis representation rather than solving for the coefficients over a fixed basis. We first show that the RWG basis is essentially a static and piecewise-linear realization of the Kolmogorov-Arnold representation theorem. Inspired by this insight, we propose PhyKAN, a physics-informed Kolmogorov-Arnold Network (KAN) that generalizes RWG into a learnable and adaptive basis family. Derived from the EFIE, PhyKAN integrates a local KAN branch with a global branch embedded with Green’s function priors to preserve physical consistency. It is demonstrated that, across canonical geometries, PhyKAN achieves sub-$10^{-2}$ reconstruction errors as well as accurate, unsupervised radar cross section predictions, offering an interpretable, physics-consistent bridge between classical solvers and modern neural network models for electromagnetic modeling.
@@ I Introduction
Solving electromagnetic (EM) scattering via the Electric Field Integral Equation (EFIE) is central to computational electromagnetics (CEM) [1], with the Method of Moments (MoM) being the standard approach. This method expands the induced surface current using predefined basis functions, typically the Rao–Wilton–Glisson (RWG) basis [2]. However, these basis functions are fixed by mesh geometry and are inherently static, a fact that prevents them from adapting to the spatial complexity of current distributions. This rigidity reveals a key limitation of MoM, as its basis functions are not learnable and cannot adapt to the underlying physics.
To address the latter limitation, several directions have been explored. Higher-order basis functions enrich MoM expressiveness using higher-degree polynomials [3, 4, 5], while characteristic or macro-basis constructions improve efficiency through physics-informed domain decomposition [6, 7]. On another front, instead of improving the basis itself, Adaptive Mesh Refinement (AMR) refines the discretization in regions of high error [8, 9]. Although effective, these approaches remain largely hand-crafted, problem-specific, and fundamentally tied to predefined basis formulations.
@@
In this letter, we reformulate MoM basis construction as a learnable function problem and show that KAN provides a theoretically grounded generalization of the RWG basis. In addition, a physics-guided EFIE-based framework, termed as PhyKAN, that embeds Green’s function [22] priors for physically consistent and adaptive current modeling, while remaining compact and interpretable, is introduced. The proposed learning framework accommodates multiple kernel functions and structural extensions, offering a unified bridge between classical integral equation solvers and neural network models.
@@ II Surface Current as a Learning Problem
Figure 1: (a) The architecture of the proposed PhyKAN comprising a local branch that encodes geometric features using a conventional KAN and global branch that incorporates Green’s function priors for physically consistent prediction of the induced surface current. (b) The structure of the conventional KAN [17], where learnable activation functions enable adaptive basis modeling.
Let $\mathbf{G}\left(\cdot\right)\in\mathbb{C}^{3\times 3}$ denote the dyadic Green’s function in the three-dimensional free space, $\boldsymbol{J}(\cdot)\in\mathbb{C}^{3\times 1}$ the induced surface current, and $\boldsymbol{E}_{\rm inc}(\cdot)\in\mathbb{C}^{3\times 1}$ the incident field. By applying the boundary conditions of Perfectly Electrically Conducting (PEC) objects $\mathcal{D}$, the EFIE can be formulated $\forall\boldsymbol{p}\in\mathbb{R}^{3}$ as:
$\int_{\boldsymbol{p}^{\prime}\in\mathcal{D}}\mathbf{G}\left(\boldsymbol{p},\boldsymbol{p}^{\prime}\right)\boldsymbol{J}\left(\boldsymbol{p}^{\prime}\right)d\boldsymbol{p}^{\prime}=\boldsymbol{E}_{\rm inc}(\boldsymbol{p}).$ (1)
The MoM discretizes this integral equation by expanding the unknown current $\boldsymbol{J}(\boldsymbol{p})$ using a set of predefined basis functions. Typically, the RWG basis $\{{f}_{n}(\cdot)\}_{n=1}^{3}$ comprising fixed and geometry-defined functions [2] is used, yielding the following surface current approximation:
@@
where $I_{n,k}$ are the unknown complex coefficients of the $n$-th RWG basis function on the $k$-th triangular element centered at $\boldsymbol{p}_{k}^{\prime}$. This geometry-fixed basis restricts the MoM’s adaptability to complex current distributions. Existing refinements, such as higher-order bases [3, 4, 5], the characteristic basis function method (CBFM) [7], and AMR[8, 9], partially alleviate this limitation but do not alter the static nature of the basis itself.
Although recent deep learning works [10, 11, 12, 13, 14, 15, 16] attempt to either learn geometry-to-current mappings or accelerate MoM via surrogate models, they still operate on fixed basis functions, thus leaving the aforedescribed core limitation unresolved. This reveals a fundamental gap: a formulation that preserves the MoM structure while enabling adaptive basis functions is still missing. To this end, instead of solving for $I_{n,k}$ in (2) satisfying the EFIE in (1), we focus on learning an optimal, data-driven basis family $\left\{b_{n}(\cdot;\boldsymbol{u})\right\}$, yielding the following generalized expansion:
@@
The latter problem finds its direct mathematical foundation in the Kolmogorov–Arnold representation theorem, which states that any multivariate function can be decomposed into finite sums of univariate functions [23]:
$f\left(x_{1},\ldots,x_{d}\right)=\sum_{i=1}^{2d+1}\alpha_{i}\left(\sum_{j=1}^{d}\phi_{ij}\left(x_{j}\right)\right),$ (4)
where $d$ is the input dimension, $\alpha_{i}$ are learnable weights, and $\phi_{ij}(\cdot)$’s are univariate nonlinear mappings. The KAN is the neural network architecture that operationalizes this theorem. To construct learnable basis functions, we embed the KAN parameterization into each $b_{n}(\cdot;\boldsymbol{u})$ as
$b_{n}\left(\boldsymbol{p}^{\prime}_{k};\boldsymbol{u}_{n}\right)=\sum_{i=1}^{M}c_{n,i}\phi\left(\boldsymbol{w}_{n,i}^{\top}\mathbf{p}_{k}^{\prime}+\beta_{n,i}\right),$ (5)
where $\phi(\cdot)$ plays the same role as the univariate nonlinear mappings $\phi_{ij}(\cdot)$ in (4), and $\left\{\boldsymbol{w}_{n,i},\beta_{n,i},c_{n,i}\right\}$ are learnable parameters controlling each kernel’s orientation, offset, and amplitude, respectively. The term $\phi\left(\boldsymbol{w}_{n,i}^{\top}\mathbf{p}_{k}^{\prime}+\beta_{n,i}\right)$ performs a learnable linear projection of the spatial coordinate $\boldsymbol{p}^{\prime}_{k}$, allowing each kernel to adapt to different geometric directions and field variations.
@@
$\boldsymbol{J}(\boldsymbol{p}^{\prime}_{k})\approx\sum_{n=1}^{N_{b}}\sum_{i=1}^{M}\alpha_{n,k}c_{n,i}\phi\left(\boldsymbol{w}_{n,i}^{\top}\mathbf{p}_{k}^{\prime}+\beta_{n,i}\right).$ (6)
To more closely mirror the MoM form in (2), we next absorb the coefficient into the spline weights via the definition $\tilde{c}_{n,i,k}\triangleq\alpha_{n,k}c_{n,i}$. This results in the following compact, MoM-consistent expansion for the surface current:
$\boldsymbol{J}(\boldsymbol{p}^{\prime}_{k})\approx\sum_{n=1}^{N_{b}}\sum_{i=1}^{M}\tilde{c}_{n,i,k}\phi\left(\boldsymbol{w}_{n,i}^{\top}\mathbf{p}_{k}^{\prime}+\beta_{n,i}\right).$ (7)
This expression is structurally isomorphic to the MoM formulation in (2), with each $\tilde{c}_{n,i,k}$ acting as the learnable counterpart of each $I_{n,k}$ and $\phi\left(\boldsymbol{w}_{n,i}^{\top}\mathbf{p}_{k}^{\prime}+\beta_{n,i}\right)$’s serving as data-driven, adaptive generalizations of the fixed RWG basis functions. Critically, unlike MoM, which only solves for $I_{n,k}$’s against a fixed basis $\{{f}_{n}(\cdot)\}_{n=1}^{3}$, we regard KAN as a learnable generalization of the RWG basis that is deployed to preserve the MoM structure while enabling physics-adaptive basis functions.
@@ III Proposed PhyKAN Framework
Figure 2: Reconstruction results for the surfaces of two objects with different variants of the proposed two-branch PhyKAN architecture.
In this section, we build upon the preceding formulation that recasts the MoM basis functions into a learnable KAN-based representation, and introduce PhyKAN, a physics-guided learning framework derived from the EFIE. The proposed artificial neural network, whose architecture is visualized in Fig. 1(a), has a dual-branch structure that enables adaptive basis learning by separately modeling local geometric features and global EM interactions.
@@ III-A Neural Network Architecture
The local branch learns edge-wise basis representations by treating each RWG basis function as a node consistent with MoM, where an RWG basis is defined on the shared edge of two triangular faces. The input is a set of $M$ edge-nodes, each described by a geometric feature vector $\mathbf{g}_{m}\in\mathbb{R}^{d}$ encoding local geometry, where $d$ is the dimension of the feature vector that encodes local geometric information. As shown in Fig. 1(a), this descriptor is then mapped through a KAN block to produce a latent feature vector, as follows:
$\mathbf{h}^{(l)}_{m}\triangleq\mathrm{KAN}_{\boldsymbol{u}}\left(\mathbf{g}_{m}\right).$ (8)
where $\mathbf{h}^{(l)}_{m}\in\mathbb{R}^{D}$ denotes the $l$-th layer embedding. As shown in Fig. 1(b)), each KAN layer uses spline-based learnable activations (with optional $B$-spline or Fourier kernels) to enable adaptive basis shaping. The outputs of the three layers are fused to form the complete latent representation of each edge embedding $\mathbf{h}_{m}$, which serves as the learned local counterpart of the conventional RWG basis.
The local feature $\mathbf{H}_{k}$ of each $k$-th element-node is aggregated from its three associated edge local embeddings of the corresponding edges ($\mathbf{h}_{a}$, $\mathbf{h}_{b}$, and $\mathbf{h}_{c}$), as follows:
@@
The global branch of PhyKAN models long-range electromagnetic interactions across the surface. In contrast to the local branch where each node corresponds to a mesh edge, here one element-node is defined per triangular element. The input to this branch is a set of $K$ element-nodes, where each node $k$ is represented by the centroid $\boldsymbol{p}_{k}\in\mathbb{R}^{3}$ of the $k$-th element, together with the incident direction $(\theta,\phi)$ of the scattered EM fields $\boldsymbol{E}_{\rm inc}(\boldsymbol{p}_{k})$’s. To incorporate EM priors into this branch, we initialize the weights of its first layer, $\mathbf{W}_{k,j}^{(0)}\in\mathbb{R}^{K\times K}$, using the dyadic Green’s function, as follows:
@@
which encodes the physical interaction between element-nodes $k$ and $j$. To incorporate the dependence on incidence, the incident angles $(\theta,\phi)$ are mapped into a $D_{\rm ang}$-dimensional (with $D_{\rm ang}\triangleq 4L$) Fourier embedding, as follows:
@@ III-B Physics-Informed Loss Function
Instead of directly minimizing the difference between predicted and reference currents, we adopt a residual-based loss function derived directly from the EFIE, which is defined as:
${L}=\frac{1}{N_{s}}\sum_{u=1}^{N_{s}}\left\|\sum_{j=1}^{K}\mathbf{G}(\boldsymbol{p}_{k},\boldsymbol{p}_{j})\boldsymbol{J}_{\rm pre}(\boldsymbol{p}_{k})-\boldsymbol{E}_{\rm inc}(\boldsymbol{p}_{k})\right\|_{2}^{2},$ (14)
where $N_{s}$ is the batch size and $\boldsymbol{p}_{k}$ is the center of the $k$-th triangular mesh element. This loss enforces the predicted surface currents to satisfy the EFIE, thereby ensuring physical consistency without requiring any supervised labels.

### v3

@@ Abstract
This letter proposes PhyKAN, a novel electromagnetics-informed learning framework that embeds the governing Electric Field Integral Equation (EFIE) directly into a dual-branch architecture for surface-current modeling. Inspired by the Kolmogorov-Arnold representation theorem, PhyKAN formulates a physics-informed framework where the local branch learns geometry-adaptive basis functions via Kolmogorov–Arnold Networks (KAN), while the global branch incorporates Green’s function priors to preserve long-range interactions and physical consistency. This design offers a unified formulation that generalizes classical Rao-Wilton-Glisson (RWG) basis into a data-driven, adaptive, and interpretable representation. It is demonstrated that, across canonical geometries, PhyKAN achieves reconstruction relative errors below 3%, preserves accurate RCS patterns, and remains robust across ablation, frequency, and mesh-resolution tests, with interpretable KAN-based complexity.
@@ I Introduction
Solving electromagnetic (EM) scattering via the Electric Field Integral Equation (EFIE) is central to computational electromagnetics (CEM) [1], with the Method of Moments (MoM) being the standard approach. This method expands the induced surface current using predefined basis functions, typically the Rao–Wilton–Glisson (RWG) basis [2]. Despite its success, MoM remains fundamentally non-learnable because both the mesh-prescribed basis space and the Green’s function operator kernel [22] are fixed, limiting its ability to adapt to geometric or physical complexity.
To address the latter limitation, several directions have been explored. Higher-order basis functions enrich MoM expressiveness using higher-degree polynomials [3, 4, 5], while characteristic or macro-basis constructions improve efficiency through physics-informed domain decomposition [6, 7]. On another front, instead of improving the basis itself, Adaptive Mesh Refinement (AMR) refines the discretization in regions of high error [8, 9]. Although effective, these approaches remain hand-crafted and problem-specific, and they retain the fixed discretization structure of MoM, leaving the overall formulation fundamentally non-adaptive.
@@
In this letter, we introduce PhyKAN, a dual-branch, physics-informed framework that embeds the EFIE directly into a learnable architecture. The key innovation is to replace the fixed Rao–Wilton–Glisson basis with KAN-based adaptive basis functions, allowing the surface-current representation to follow geometric and physical variations. Meanwhile, the global branch incorporates the Green’s function kernel as an explicit prior, lifting it into a higher-dimensional latent space to model long-range coupling while preserving EFIE consistency. By uniting KAN-driven basis adaptation with a Green’s-function-embedded global prior, PhyKAN offers a compact, physically grounded reformulation of EFIE–MoM within a modern neural architecture.
@@ II Surface Current as a Learning Problem
Let $\mathbf{G}(\cdot)\in\mathbb{C}^{3\times 3}$ denote the dyadic Green’s function in three-dimensional free space, $\boldsymbol{J}(\cdot)\in\mathbb{C}^{3\times 1}$ the induced surface current, and $\boldsymbol{E}_{\rm inc}(\cdot)\in\mathbb{C}^{3\times 1}$ the incident field. Let $\boldsymbol{n}(\boldsymbol{p})$ denote the outward unit normal vector at $\boldsymbol{p}$ on the PEC surface $\mathcal{S}$. By applying the boundary conditions of a Perfectly Electrically Conducting (PEC) object with surface $\mathcal{S}$, the EFIE can be formulated $\forall\boldsymbol{p}\in\mathcal{S}$ as:
$\mathbf{n}(\boldsymbol{p})\times\left[\mathbf{E}_{\text{inc }}(\boldsymbol{p})+\int_{\mathcal{S}}\mathbf{G}\left(\boldsymbol{p},\boldsymbol{p}^{\prime}\right)\boldsymbol{J}\left(\boldsymbol{p}^{\prime}\right)\mathrm{d}\mathcal{S}^{\prime}\right]=\mathbf{0},\quad\boldsymbol{p}\in\mathcal{S}$ (1)
Here, $\mathbf{G}\left(\boldsymbol{p},\boldsymbol{p}^{\prime}\right)$ governs the electromagnetic interaction, while $\boldsymbol{J}\left(\boldsymbol{p}^{\prime}\right)$ is expanded through MoM basis functions. Together, these two elements define the core structure that PhyKAN reformulates into a learnable and physically consistent representation.
II-A Learning the Basis
The MoM discretizes this integral equation (1) by expanding the unknown current $\boldsymbol{J}(\boldsymbol{p})$ using a set of predefined basis functions. Typically, the RWG basis $\{{f}_{n}(\cdot)\}_{n=1}^{3}$ comprising fixed and geometry-defined functions [2] is used, yielding the following surface current approximation:
@@ II-A Learning the Basis
where $I_{n,k}$ are the unknown complex coefficients of the $n$-th RWG basis function on the $k$-th triangular element centered at $\boldsymbol{p}_{k}^{\prime}$. This geometry-fixed basis restricts the MoM’s adaptability to complex current distributions.
In the classical MoM formulation, all degrees of freedom reside in the scalar coefficients $\{I_{n,k}\}$, whereas the basis shapes $\{f_{n}(\cdot)\}$ are entirely dictated by the mesh geometry and remain identical across different illuminations or target configurations. The RWG family is edge-based and divergence-conforming, which makes it physically sound but also rigid: improving expressiveness typically relies on mesh refinement or higher-order/macrobasis constructions, increasing the number of unknowns rather than enriching the basis itself.
To introduce adaptability while preserving the MoM structure, we reinterpret basis construction itself as a learnable function approximation problem. Rather than only solving for fixed coefficients $I_{n,k}$ in (2) that scale a geometry-defined basis, we seek an optimal, data-driven basis family $\{b_{n}(\cdot;\boldsymbol{u})\}$ whose functional shape can adapt to the local geometry and electromagnetic response, while the expansion in (3) keeps the familiar MoM form, yielding the following generalized expansion:
@@
This formulation is enabled by the Kolmogorov–Arnold representation theorem [23], which expresses multivariate functions as compositions of univariate nonlinear mappings. Motivated by this structure, each basis function is parameterized using a KAN representation:
$b_{n}\left(\boldsymbol{p}^{\prime}_{k};\boldsymbol{u}_{n}\right)=\sum_{i=1}^{M}c_{n,i}\phi\left(\boldsymbol{w}_{n,i}^{\top}\boldsymbol{p}_{k}^{\prime}+\beta_{n,i}\right),$ (4)
where $\phi(\cdot)$ plays the same role as the univariate nonlinear mappings $\phi_{ij}(\cdot)$ in [23], and $\left\{\boldsymbol{w}_{n,i},\beta_{n,i},c_{n,i}\right\}$ are learnable parameters controlling each kernel’s orientation, offset, and amplitude, respectively. The term $\phi\left(\boldsymbol{w}_{n,i}^{\top}\boldsymbol{p}_{k}^{\prime}+\beta_{n,i}\right)$ performs a learnable linear projection of the spatial coordinate $\boldsymbol{p}^{\prime}_{k}$, allowing each kernel to adapt to different geometric directions and field variations.
@@
$\boldsymbol{J}(\boldsymbol{p}^{\prime}_{k})\approx\sum_{n=1}^{N_{b}}\sum_{i=1}^{M}\tilde{c}_{n,i,k}\phi\left(\boldsymbol{w}_{n,i}^{\top}\boldsymbol{p}_{k}^{\prime}+\beta_{n,i}\right),$ (5)
with $\tilde{c}_{n,i,k}\triangleq\alpha_{n,k}c_{n,i}$ acting as the learnable counterpart of each $I_{n,k}$ and $\phi\left(\boldsymbol{w}_{n,i}^{\top}\boldsymbol{p}_{k}^{\prime}+\beta_{n,i}\right)$’s serving as data-driven, adaptive generalizations of the fixed RWG basis functions. The resulting expansion mirrors the MoM form in (2) while replacing the static RWG basis with a KAN-based, physics-adaptive representation.
II-B Embedding Green’s Function Physics
The dyadic Green’s function $\mathbf{G}\left(\boldsymbol{p},\boldsymbol{p}^{\prime}\right)$ in (1) fully characterizes the electromagnetic interaction kernel in the EFIE and governs long-range coupling across the surface. In classical MoM, this operator is fixed and directly used to assemble the system matrix, leaving no room for learning enriched representations of global physics. To introduce learnability without violating the EFIE structure, PhyKAN incorporates $\mathbf{G}$ as an explicit physics prior and then lifts this operator into a higher-dimensional latent space. Formally, for the $k$-th element we construct
$\mathbf{F}_{k}=\Phi\left(\mathbf{G}\left(\boldsymbol{p}_{k},\boldsymbol{p}_{k^{\prime}}\right)\right)$ (6)
where $\Phi(\cdot)$ is a linear mapping initialized to satisfy $\Phi(\mathbf{G})=\mathbf{G}$, ensuring exact physical fidelity at the start of training. During optimization, $\Phi$ lifts the Green’s kernel $\mathbf{G}$ into a latent representation $\mathbf{F}_{k}\in\mathbb{C}^{D}$($D>3$) preserving the EFIE prior while enabling enriched, geometry-aware representations of long-range coupling.
This high-dimensional lifting contrasts sharply with classical MoM, where the Green’s kernel is confined to a fixed $3\times 3$ dyadic structure. By embedding $\mathbf{G}$ as a prior and expanding its representational space, PhyKAN retains strict physical consistency while unlocking additional flexibility for capturing complex, nonlocal interactions.
@@ III Proposed PhyKAN Framework
Figure 1: (a) PhyKAN consists of a KAN-based local branch that encodes geometry-adaptive basis features and a global branch that incorporates Green’s function priors for physically consistent modeling of long-range electromagnetic interactions. (b) The conventional KAN structure [17], where learnable activation functions form adaptive univariate kernels.
In this section, we build upon the preceding formulation into PhyKAN, an EFIE-derived, physics-guided learning framework. The proposed artificial neural network, whose architecture is visualized in Fig. 1(a), has a dual-branch structure that enables adaptive basis learning by separately modeling local geometric features and global EM interactions.
@@ III-A Neural Network Architecture
The local branch learns edge-wise basis representations by treating each RWG basis function as a node, consistent with the MoM formulation in which each RWG function is defined on the shared edge of two triangular faces. The input is a set of $M$ edge-nodes, each described by a geometric feature vector $\mathbf{g}_{m}\in\mathbb{R}^{d}$ where the dimension $d$ of the feature vector includes edge length, adjacent-face normals, and relative edge orientation. As illustrated in Fig. 1(a), the geometric descriptor is passed through a three-layer KAN encoder. Each KAN layer performs a nonlinear lifting of $\mathbf{g}_{m}$ through spline-based learnable activations:
$\mathbf{h}^{(l)}_{m}\triangleq\mathrm{KAN}_{\boldsymbol{u}}\left(\mathbf{g}_{m}\right),$ (7)
where $\mathbf{h}^{(l)}_{m}\in\mathbb{R}^{D}$ denotes the $l$-th layer embedding. Each layer contains learnable linear projections and kernel activations (B-spline, rational, or Fourier), enabling the basis to adapt its local functional shape.
To obtain a stable and expressive embedding, the three KAN outputs are fused via a residual linear aggregator:
$\mathbf{h}_{m}=\sum_{l=1}^{3}{\eta}_{l}\odot\mathbf{h}_{m}^{(l)},$ (8)
where ${\eta}_{l}$ are learnable per-layer weighting vectors that fuse multi-level KAN features into the final geometry-adaptive edge embedding.
The local feature $\mathbf{H}_{k}$ of each $k$-th triangular element is aggregated from its three associated edge local embeddings of the corresponding edges ($\mathbf{h}_{a}$, $\mathbf{h}_{b}$, and $\mathbf{h}_{c}$), as follows:
@@
The global branch of PhyKAN models long-range electromagnetic interactions across the surface. In contrast to the local branch where each node corresponds to a mesh edge, here one element-node is defined per triangular element. The input to this branch is a set of $K$ element-nodes, where each node $k$ is represented by the centroid $\boldsymbol{p}_{k}\in\mathbb{R}^{3}$ of the $k$-th element, together with the incident direction $(\theta,\phi)$ of the scattered EM fields $\boldsymbol{E}_{\rm inc}(\boldsymbol{p}_{k})$’s. To embed electromagnetic physics into this branch, the first-layer coupling weights between element-nodes are initialized directly with the dyadic Green’s function:
@@
which encodes the physical interaction between elements $k$ and $j$ as governed by the EFIE kernel. This initialization ensures that global message passing is grounded in the true nonlocal behavior of Maxwell’s equations.
To incorporate the dependence on incidence, the incident angles $(\theta,\phi)$ are mapped into a $D_{\rm ang}$-dimensional (with $D_{\rm ang}\triangleq 4L$) Fourier embedding, as follows:
@@ III-B Physics-Informed Loss Function
For each incident field $\boldsymbol{E}^{(m)}_{\text{inc}}$, we compute the EFIE residual over all $K$ surface elements, rather than supervising the currents with MoM solutions. The residual for the $m$-th incident field is written as:
$\mathcal{L}^{(m)}=\frac{1}{K}\sum_{k=1}^{K}\left\|\boldsymbol{n}(\boldsymbol{p}_{k})\times\left(\boldsymbol{E}^{(m)}_{\text{inc}}(\boldsymbol{p}_{k})+\sum_{j=1}^{K}\boldsymbol{G}(\boldsymbol{p}_{k},\boldsymbol{p}_{j})\boldsymbol{J}^{(m)}_{\text{pre}}(\boldsymbol{p}_{j})\right)\right\|_{2}^{2}$ (14)
where $\boldsymbol{p}_{k}$, $\boldsymbol{p}_{j}$ denote the centers of the centers of the $k$-th and $j$-th mesh elements. This residual enforces the predicted current $\boldsymbol{J}^{(m)}_{\text{pre}}$ to satisfy the PEC boundary condition for the incident field $\boldsymbol{E}^{(m)}_{\text{inc}}$. The overall physics-informed loss is obtained by averaging over a batch of $N_{s}$ incident fields:
$\mathcal{L}_{\text{EFIE}}=\frac{1}{N_{s}}\sum_{m=1}^{N_{s}}\mathcal{L}^{(m)}.$ (15)
This formulation enables PhyKAN to be trained without MoM current labels, relying purely on the discretized EFIE for physical supervision.